OpenStudy (anonymous):

Help please! Solve the following system of equations: -2x + y = 1 -4x + y = -1 (3,1) (-1,3) (-1,-3) (1,3)

OpenStudy (compassionate):

-2x + y = 1 -4x + y = -1 Multiply by -1 and add -2x + y = 1 4x - y = 1 2x = 2 x = 1 Now plug in x to solve for y.

OpenStudy (anonymous):

I don't understand @Compassionate